Tax incidence refers to who “feels” the effects of a tax, buyers or sellers. On the example from the previous page, the $4 tax resulted in buyers paying $3 more (since the market price was $4 without the tax, and the price buyers pay with the tax is $7) and sellers receiving $1 less (since the market price was $4 without the tax and sellers receive $3 with the tax). So, we say the incidence of the $4 tax is $3 on buyers and $1 on sellers. Note that the tax incidence is the same whether the tax is physically paid by buyers or sellers. This is important – making sellers pay the tax won’t help or hurt consumers. The effect of the tax is the same no matter who pays. Suppose before a tax is created, the market price is $10.09 and the quantity is 120. Now suppose a government creates a $5.6 excise tax paid by sellers. After the tax, the quantity falls to 100 and the market price is $12.12. Based on this information, the tax incidence on sellers is:
Answer #1

Tax incidence on buyer=12.12-10.09=2.03

tax=5.6

tax incidence on seller=5.6-2.03=3.57